Circular No. 8138

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 K2 It has been noted by numerous individuals that the preliminary orbital

elements of comet C/2003 K2 (cf. IAUC 8136) place it close to the position of an unconfirmed object found on SWAN ultraviolet SOHO website images and reported to the Central Bureau on Apr. 14 by X.-m. Zhou (Bo-le, Xin-jiang, China). Measurements of the object on six dates, Apr. 5–19, were forwarded to the Central Bureau by Zhou (via D. H. Chen), by M. Mattiazzo, and by S. Hönig; the positions differed considerably, due to the poor resolution of SWAN (uncertainty on the order of 1^{o}). Two search ephemerides based on various positions were circulated by the Bureau to numerous visual and CCD observers in the hopes of optical confirmation, but the searches (undertaken during the last week of April by Zhou, A. Hale, Mattiazzo, Y. Kushida, and Y. Ezaki) revealed nothing to as faint as mag 14.5. The following improved parabolic orbital elements for C/2003 K2 (from MPEC 2003-K49) indicate that the search-ephemeris positions in late April for the SWAN object were no closer than ∼ 2^{o}.5 from C/2003 K2. The comet might be of short period.

T = 2003 Apr. 10.550 TT  ω = 357.^{o}591
Ω = 89.942 }2000.0
q = 0.62325 AU  i = 10.273

COMET C/2003 K3 (SOHO) Further to IAUC 8120, D. Hammer reports measurements for a faint

comet (mag 8–10) found by H. Otterstedt on SOHO website C3 images. The reduced astrometry, orbital elements (ω = 108^{o}, Ω = 79^{o}, i = 39^{o}, q = 0.14 AU), and ephemeris by B. G. Marsden appear on MPEC 2003-K46.

2003 UT  α_{2000}  δ_{2000}
May 25.096  4^{h}29.^{m}3  +23^{o}24′

S/2003 J 21 Further to IAUC 8125, the discovery of a new satellite of Jupiter has

been reported on CCD images taken with the 3.6-m Canada-France-Hawaii Telescope by teams led by B. Gladman and by S. S. Sheppard. Astrometry, preliminary orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2003-K45 (e = 0.21, i = 148^{o}, P = 600 days, H = 16.3).

2003 May 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green